OpenStudy (anonymous):

Question from my book. Prove that \(n^2 > 7n +1\) for all \(n \geq 8 \). The result is true when \(n=8\) because \(8^2 = 64\) and \(7x8 +1 = 57\). Suppose it is true when \(n\) is any number \(k \geq 8\), that is \(k^2 > 7k+1\). Then \[ (k+1)^2 = k^2 +2k +1 > (7k+1) +2k+1 = 7(k+1)+1+(2k-6)\] Since \(k \geq 8\), \(2k-6\) is a natural number and the last expression above is greater than \(7(k+1)+1\). The induction step is verified, and so the result is true for all \(n \geq 8\). Could someone explain this for me? How is this a valid proof? It is supposed to be a about induction.

OpenStudy (mr.math):

Read this http://cms.dt.uh.edu/faculty/delavinae/F05/2305/Induction.PDF

OpenStudy (slaaibak):

You first have to proof for n=8, the base case. Now if this is true, you need to assume it's true for n. Then you have to prove it for n+1 When you prove it for n+1, you prove it for 9, then 9 becomes n, so you prove it for 9+1=10 This process goes on until infinity. Therefore it proves for all n>=8

OpenStudy (mr.math):

I think his question is why the principle of mathematical induction is true. In the link I gave above you can find the proof of the principle theorem.

OpenStudy (anonymous):

I don't argue with induction. just with this problem, how does this line make sense "Since k≥8, 2k−6 is a natural number and the last expression above is greater than 7(k+1)+1. "

OpenStudy (anonymous):

@Mr.Math reading it now

OpenStudy (zarkon):

2*k-6>0 for k>=8

OpenStudy (mr.math):

Oh so you're just talking about this problem, sorry!

OpenStudy (zarkon):

so \[ 7(k+1)+1+(2k-6)\ge 7(k+1)+1\]

OpenStudy (slaaibak):

I would have done the proof a bit different. \[k^2 + 2k + 1 > 7(k + 1) +1\] \[(k^2) + (2k +1) > (7k + 1) + (7)\] we know k^2 > 7k + 1 and we know 2k + 1 > 7 since k>=8

OpenStudy (zarkon):

I would have used calculus ;)

OpenStudy (anonymous):

@Zarkon I would too.

OpenStudy (anonymous):

Can we agree that the example given above is not good?